A machine can fill 3864 cartons of milk in 0.4 hour. How many cartons of milk can it fill per hour?
I came up this 7728. Can you tell me if this is correct?

Answer by stanbon(57224)

A machine can fill 3864 cartons of milk in 0.4 hour. How many cartons of milk can it fill per hour?
I came up this 7728. Can you tell me if this is correct?
-----------------
3864/0.4 = x/1
cross-multiply 0.4x = 3864
x = 3864/0.4
x = 9660 cartons in 1 hour.
=============
Cheers,
Stan H.